write an equation of the line through the point (4, -2) that is parallel to the line -5x + 6y = -2
Answer by Alan3354(69443)

write an equation of the line through the point (4,-2) that is parallel to the line -5x + 6y = -2
---------
Step 1, find the slope of the line, m
Step 2, use y-y1 = m*(x-x1) where (x1,y1) is the point.